OpenStudy (anonymous):

Use the information for the isotopes of element X to calculate the average atomic mass of X, and identify the element. Isotope Abundance Mass (amu) 107X 51.85% 106.90509 109X 48.15% 108.90470 Help is much appreciated, a bit confused? Could you calculate and explain how? Thank much.

OpenStudy (aaronq):

weighted average (mass 1 x abundance of 1)+ (mass 2 x abundance of 2)= sum of both masses

OpenStudy (anonymous):

Now that I have the sum of both masses what do i do?

OpenStudy (aaronq):

sorry where i wrote sum of both masses, i mean weighed sum of masses, meaning average atomic mass

OpenStudy (anonymous):

Okay so I have the sum of both masses, how do I proceed from here? I'm still confused haha.

OpenStudy (aaronq):

(mass 1 x abundance of 1)+ (mass 2 x abundance of 2)= average atomic mass

OpenStudy (aaronq):

You don't just add the masses

OpenStudy (anonymous):

Yes so I have the average atomic mass, but how does this help me find the element? Is the average atomic mass the mass of the element?

OpenStudy (aaronq):

Average atomic mass is the same as molar mass, so use a periodic table

OpenStudy (anonymous):

106.9 *51.85 + 108.9*48.15 = 12887.3475 Okay so I have this, what did I do wrong? This is obviously not present on the periodic table.

OpenStudy (aaronq):

you didn't use the percentages correctly. 51.85%= 0.5185

OpenStudy (anonymous):

Thanks so much for your help!

OpenStudy (aaronq):

no problem !